Suppose there are two groups of people in an economy, with their respective consumption function.

                        CA=100+0.5YA

                        CB=150+0.75YB

   planned investment 200.Where Ci and Yare respectively consumption and income level

 for the ith group (for all i=A,B). Suppose each group gets 50% of total GDP in the economy.

Determine the equilibrium level of income. Considering the equilibrium income as the initial income, a lumpsum tax T is imposed on Group A and the same amount is given as a transfer to Group B.

 Now determine planned consumption, planned savings , planned aggregate Demand and the Actual Investment as the initial level of income



1
Expert's answer
2021-06-17T09:47:58-0400

The equilibrium level of income is:

Y = Ca + Cb + I = 100 + 0.5×0.5Y + 150 + 0.75×0.5Y + 200 = 450 + 0.625Y,

0.375Y = 450,

Y = 1,200.

If a lumpsum tax T is imposed on Group A and the same amount is given as a transfer to Group B, then Ca will decrease, and Cb will increase.